The results below were obtained at failure in a series of consolidated-undrained (CU) triaxial tests, with pore pressure measurement, on specimens of fully saturated clay. Determine the values of the shear strength parameters cr, ??, c, and ?, using both total and effective Mohr Circles. 150 300 450 600 103 202 305 410 169 252 331 O3 (kPa) u (la)

2.If an additional CU triaxial test with pore pressure measurement is to be performed using a ?3 of 250 kPa on a specimen identical to those in last Problem , estimate the deviator stress (?d) and pore pressure (u) of this additional test at failure. What would potentially be orientation of the failure plane in the specimen?
Answer #1

(utun duta 30D 450 20s-410 2s2 31 10 3 20 2 8 2 69 2S3 U CKPa) Sdulron A$2.18 Poroln ingtshe on ふ lus ニう| sd = !ье.qgkPa . | Au lus )-(2-u) tur-se.os-ナ2 18.98-L)(2)x 2 (2184) 422.74-u = 64275-2.57 íxu 1:57( u=219.